| #1225 |
28 Nov 10:00 Log Created Nov 2016 Campaign
|
|
Start experiment
|
normal
|
|
|
| #1226 |
28 Nov 15:00 Antenna VSWR
|
|
Antennas
|
normal
|
|
|
| #1228 |
28 Nov 23:00 Transmitter Power Measurement
|
|
Transmitters
|
normal
|
|
|
| #1230 |
29 Nov 03:40 Fail TX3
|
|
Transmitters
|
normal
|
|
|
| #1231 |
29 Nov 07:00 Transmitter Power Measurements
|
|
Transmitters
|
normal
|
|
|
| #1233 |
28 Nov 17:50 Start of the experiment
|
I. Manay
|
Start experiment
|
normal
|
|
|
| #1234 |
28 Nov 19:00 Vipir system was stopped
|
I. Manay
|
Reception
|
normal
|
|
|
| #1235 |
28 Nov 19:00 Spark gaps adjustment history
|
W. Rojas
|
Spark Gaps
|
normal
|
|
|
| #1236 |
28 Nov 19:59 Operation Resume
|
I. Manay
|
General
|
normal
|
|
|
| #1237 |
29 Nov 19:30 Transmitter power measurements
|
|
Transmitters
|
normal
|
|
|
| #1238 |
29 Nov 08:12 Channel B EW Drift
|
|
General
|
normal
|
|
|
| #1239 |
29 Nov 10:03 VIPIR is turned off
|
|
General
|
normal
|
|
|
| #1240 |
29 Nov 10:45 Transmitter on
|
|
Transmitters
|
normal
|
|
|
| #1241 |
29 Nov 12:20 Digisonde off
|
|
Other Instruments
|
normal
|
|
|
| #1242 |
29 Nov 12:45 Transmitter on
|
|
Transmitters
|
normal
|
|
|
| #1243 |
29 Nov 15:33 EW Drift
|
|
General
|
normal
|
|
|
| #1244 |
29 Nov 16:05 VIPIR on
|
|
Other Instruments
|
normal
|
|
|
| #1245 |
29 Nov 18:55 Spark Gap adjustments
|
|
Spark Gaps
|
normal
|
|
|
| #1246 |
29 Nov 19:44 Noise
|
|
Reception
|
normal
|
|
|
| #1247 |
29 Nov 19:44 Operation Resume
|
|
General
|
normal
|
|
|
| #1248 |
29 Nov 23:00 Transmitter Power Measurements
|
|
Transmitters
|
normal
|
|
|
| #1249 |
29 Nov 23:59 Plots
|
|
Plot
|
normal
|
Others
|
|
| #1250 |
29 Nov 21:09 Pc ARES had hanged up
|
avaldez
|
Acquisition system
|
normal
|
|
|
| #1251 |
30 Nov 06:00 Spark gaps adjustment
|
A Huamani
|
Spark Gaps
|
normal
|
|
|
| #1252 |
30 Nov 09:43 Restart pc ARES
|
P.Condor
|
Acquisition system
|
normal
|
|
|
| #1253 |
30 Nov 11:35 Now is running new programs in Faraday
|
P.Condor
|
Process/Analysis Program
|
normal
|
|
|
| #1254 |
30 Nov 18:07 Spark gaps adjustment
|
A. Coronado
|
Spark Gaps
|
normal
|
|
|
| #1255 |
30 Nov 18:02 Starting Imaging
|
P.Condor
|
Acquisition system
|
normal
|
|
|
| #1256 |
30 Nov 19:03 Noise levels
|
P.Condor
|
General
|
normal
|
|
|
| #1257 |
30 Nov 19:10 Imaging off
|
P.Condor
|
Acquisition system
|
normal
|
|
|
| #1258 |
30 Nov 19:30 Transmitter power measurements
|
|
Transmitters
|
normal
|
|
|
| #1259 |
30 Nov 20:00 Figures
|
P. Condor
|
Process/Analysis Program
|
normal
|
|
|
| #1260 |
30 Nov 19:59 Operation Resume
|
P.Condor
|
General
|
normal
|
|
|
| #1261 |
30 Nov 23:55 Plots of the day
|
I. Manay
|
Plot
|
normal
|
|
|
| #1263 |
01 Dec 06:18 Spark gaps adjustment history
|
J. Lavado
|
Spark Gaps
|
normal
|
|
|
| #1264 |
01 Dec 05:15 Imaging experiment on / off
|
I. Manay
|
General
|
normal
|
|
|
| #1265 |
01 Dec 06:00 Restart of PC ARES
|
I. Manay
|
Acquisition system
|
normal
|
|
|
| #1266 |
01 Dec 05 :20 Noise history
|
I. Manay
|
Reception
|
normal
|
|
|
| #1267 |
01 Dec 07:59 Operation Resume
|
I. Manay
|
General
|
normal
|
|
|
| #1268 |
01 Dec 19:00 Transmitter Power Measurements
|
|
Transmitters
|
normal
|
|
|
| #1273 |
02 Dec 00:10 Figures
|
P. Condor
|
Process/Analysis Program
|
normal
|
|
|
| #1274 |
02 Dec 06:17 Imaging is off
|
P.Condor
|
Acquisition system
|
normal
|
|
|
| #1276 |
02 Dec 06:22 Restart Faraday because was desynchronized
|
P.Condor
|
Acquisition system
|
normal
|
|
|
| #1277 |
02 Dec 06:19 Noise levels
|
P.Condor
|
General
|
normal
|
|
|
| #1278 |
02 Dec 06:03 Spark gaps adjustment
|
W. Rojas
|
Spark Gaps
|
normal
|
|
|
| #1279 |
02 Dec 07:55 Figures
|
P. Condor
|
Process/Analysis Program
|
normal
|
|
|
| #1280 |
02 Dec 08:00 Operation Resume
|
P.Condor
|
General
|
normal
|
|
|
| #1281 |
02 Dic 19:30 Transmitter power measurements
|
|
Transmitters
|
normal
|
|
|
| #1282 |
02 Dec 23:00 Transmitter Power Measurements
|
|
Transmitters
|
normal
|
|
|
| #1285 |
03 Dec 06:12 Imaging is off
|
abelleza
|
End experiment
|
normal
|
|
|
| #1286 |
03 Dec 06:21 Spark gaps adjustment
|
abelleza
|
Spark Gaps
|
normal
|
|
|
| #1287 |
03 Dec 08:00 Transmitter Power Measurements
|
|
Transmitters
|
normal
|
|
|
| #1288 |
02 Dec 11:11 Restart Faraday
|
J.Espinoza
|
Acquisition system
|
normal
|
|
|
| #1289 |
02 Dec 17:25 Start Imaging
|
J.Espinoza
|
Acquisition system
|
normal
|
|
|
| #1290 |
02 Dec 16:00 Restart Faraday
|
J.Espinoza
|
Acquisition system
|
normal
|
|
|
| #1291 |
02 Dec 20:00 Operation Resume
|
J.Espinoza
|
General
|
normal
|
|
|
| #1292 |
02 Dec 18:33 Spark gaps adjustment
|
J.Espinoza
|
Spark Gaps
|
normal
|
|
|
| #1293 |
03 Dec 08:10 Figures
|
A. Belleza
|
Process/Analysis Program
|
normal
|
|
|
| #1294 |
03 Dec 08:01 Operation Resume
|
A. Belleza
|
General
|
normal
|
|
|
| #1295 |
06 Dec 12:00 Transmitters operative parameters
|
|
Transmitters
|
normal
|
|
|
| #1334 |
23 Jan 14:00 Final Report and Operativity
|
|
End experiment
|
normal
|
|
|